### v1

Abstract
The Segment Anything Model (SAM) has gained significant attention for its impressive performance in image segmentation. However, it lacks proficiency in referring video object segmentation (RVOS) due to the need for precise user-interactive prompts and limited understanding of different modalities, such as language and vision. This paper presents the RefSAM model, which for the first time explores the potential of SAM for RVOS by incorporating multi-view information from diverse modalities and successive frames at different timestamps. Our proposed approach adapts the original SAM model to enhance cross-modality learning by employing a lightweight Cross-Modal MLP that projects the text embedding of the referring expression into sparse and dense embeddings, serving as user-interactive prompts. Subsequently, a parameter-efficient tuning strategy is employed to effectively align and fuse the language and vision features. Through comprehensive ablation studies, we demonstrate the practical and effective design choices of our strategy. Extensive experiments conducted on Ref-Youtu-VOS and Ref-DAVIS17 datasets validate the superiority and effectiveness of our RefSAM model over existing methods. The code and models will be made publicly at github.com/LancasterLi/RefSAM.
keywords:
[…]
1 Introduction
Referring video object segmentation (RVOS) aims to segment the target object in the video with the guidance of given language expressions [1, 2]. It has attracted widespread attentions in the computer vision community due to its great potential in real world applications, such as video retrieval and video editing [3, 4, 5]. It is a challenging task and the key lies in how to incorporate multi-view information from video frames at different timestamps and align sources from different modalities, i.e., vision and language views. Compared with video object segmentation [6, 7], RVOS is more challenging as it is required to perform semantic-level alignment between vision and language modalities without the ground-truth mask annotation in the first frame. Although many efforts have been devoted to tackling this task [8, 9, 10, 11], it still remains a challenge how to accurately identify and segment the target object in the video by fully exploiting the semantic meanings of sources from vision and language views.
Recently, the Segment Anything Model (SAM) [12] has been proposed, which serves as a foundation model for image segmentation. Based on the designed prompt engineering, it can be transferred to the new task in a zero-shot manner. Due to its impressive performance on image segmentation, many works have explored to apply SAM in other related fields [13, 14], e.g., medical image analysis [15, 16, 17], remote sensing images [18, 19, 20], video object tracking [21, 22], style transfer [23], and 3D reconstruction [24]. Although these methods have achieved great progress on these related tasks based on SAM, none of them can be applied to RVOS directly due to the inherent complexity of this task. Concretely, RVOS requires the model to have the comprehensive understanding of sources from vision and language modalities. Unlike video object segmentation (VOS), no ground-truth mask annotations in the first frame are available in RVOS. It’s worth noting that although some SAM-based multiple object tracking models such as SAM-Track [21] and TAM [22] can also be adapted to the task of RVOS by providing bounding boxes of the first frame detected by extra object detection models, such as Grounding DINO [25], some non-negligible issues can exist in such frameworks. For example, they cannot work effectively in an end-to-end manner due to the two-stage pipeline. As a result, inaccurate bounding boxes in the first frame detected by the object detection model can result in the poor performance of the downstream segmentation task. On the other hand, the involvement of separate models can bring extra difficulties in the model training and deployment process.
Thus, a natural question arises that how to effectively adapt SAM to the RVOS task in an end-to-end manner to fully unleash its potential capacity for segmentation. In this paper, we conduct the initial exploration and propose the RefSAM, which is the first end-to-end SAM-based framework for the task of RVOS. Based on the powerful foundation model SAM, RefSAM can perform accurate target object segmentation in the video with given language expressions. Specifically, to enhance the cross-modality learning capability of the original SAM, we propose a lightweight Cross-Modal MLP that projects the text embedding of the referring expression into sparse and dense embeddings, serving as user-interactive prompts like points and bounding boxes prompts. Subsequently, a parameter-efficient tuning strategy is employed to effectively align and fuse the language and vision features.
Our contributions can be concluded as follows.
  1. $\bullet$
    We conduct the pioneering study to explore SAM’s potential for RVOS through the integration of multi-view information from diverse modalities and successive frames at different timestamps.
  2. $\bullet$
    We introduce RefSAM, a novel approach that utilizes lightweight modules and an efficient fine-tuning strategy to align and fuse language and vision features in an end-to-end learning manner, thereby effectively adapting SAM for RVOS.
  3. $\bullet$
    Extensive experiments on Ref-Youtu-VOS and Ref-DAVIS17 datasets demonstrate the promising results of RefSAM and its superiority compared to existing methods.
2 Related Work
2.1 Referring Video Object Segmentation
RVOS uses language cues to segment an object in a video. Unlike the VOS task which provides the ground truth bounding box of the first frame, it exploits a different type of guidance, i.e., language expressions, to identify and segment the object referred by the given language expression in a video. Therefore, the task is more challenging due to the great domain gap between language and vision modalities.
[26] first proposes the RVOS task and extends the A2D [27] and J-HMDB [28] datasets with sentences describing the actors and actions appearing in the video content. [29] extends language grounding models to video data to ensure temporally coherent predictions and augment Davis datasets [6, 30] with language descriptions. URVOS [31] proposes a unified end-to-end deep neural network that performs both language-based object segmentation and mask propagation in a single model and constructs a large-scale RVOS dataset Refer-Youtube-VOS. RefVOS [32] leverages the last stage feature of the vision backbone to fuse with the textual feature. ClawCraneNet [33] leverages three kinds of object-level relations to progressively construct discriminative visual embedding. CITD [2] puts forward a two-stage, top-down RVOS solution. MRSA [34] presents a multi-level visual representation framework that consists of video, frame, and object granularity. MMVT [9] incorporates the motion information from optical flow maps with appearance and linguistic features for text-based video segmentation. LBDT [8] proposes a Language-Bridged Duplex Transfer module to conduct spatial-temporal interaction explicitly between two independent 2D ConvNets in the encoding phase for RVOS. PMINet [10] performs multimodal interaction in each stage of the visual backbone, and multimodal feature is incorporated back into the visual backbone to guide the progressive learning of visual features. [11] utilizes implicitly individual word features to integrate visual features. These models still do not achieve outstanding performance due to the significant semantic gap between visual information and language information.
Vision Transformers (ViTs) [35] have attracted lots of attention recently and have been widely used in various domains [36, 37, 38, 39, 40, 41], including the field of RVOS. MTTR [1] models the RVOS task as a parallel sequence prediction problem and outputs all objects in the video prior to selecting the one referred by the text. ReferFormer [4] introduces a small set of object queries that are conditioned on the text expression to attend to the referred object only. VLT [42] utilizes word features for query generation and spatial dynamic fusion. Different from these models, we present a simple yet effective solution by adapting SAM to RVOS via some minimal designs.
2.2 Segment Anything Model
Recently, SAM [12] build a foundation model for segmentation. The model designs a promptable segmentation task and it can transfer zero-shot to new image distributions and tasks. Following that, a large number of research studies based on SAM emerged. SAM has been widely applied in various fields [13, 14], e.g., medical image analysis [15, 16, 17], remote sensing images [18, 19, 20], video object tracking [21, 22, 43, 44], style transfer [23], and 3D reconstruction [24]. In the medical image analysis domain, SAMed [45] involves customizing the SAM model for medical image segmentation by applying the low-rank-based (LoRA) fine-tuning strategy to the SAM image encoder which can perform better for semantic segmentation tasks on medical images. [46] demonstrates that SAM can achieve high segmentation accuracy for brain tumor MRI datasets in a point-to-mask setting. [47] assesses the zero-shot segmentation performance of the SAM model on representative segmentation tasks in whole slide imaging. In the remote sensing images domain, [19] introduces a domain-specific decoder, which learns the problem-specific semantics with only five labeled images. [20] develops an efficient pipeline with SAM to generate a large-scale remote sensing segmentation dataset named SAMRS. In the style transfer domain, Any-to-Any Style Transfer [48] enables users to specify which style region to select and which content regions to apply during style transfer. In the 3D reconstruction domain, SA3D [49] can segment any object in a 3D scene with one-shot manual prompting in a single rendered view. These advancements have shown the versatility of the foundation segmentation model.
PerSAM [50] proposes a training-free Personalization approach and a one-shot fine-tuning variant that only trains 2 parameters within 10 seconds for improved performance. HQ-SAM [51] designs a learnable High-Quality output token which injected into SAM’s mask decoder to predict the high-quality mask. These models all focus on the improvement of SAM in the vision field. SAM-Track [21] and TAM [22] enable users to select multiple objects in videos for tracking [52]. However, none of these models are currently available for the RVOS task. In this paper, we propose an end-to-end RVOS model based on SAM for the first time.
3 RefSAM for Reffering Video Object Segmentation
First, we present the preliminaries of SAM in Sec. 3.1. Then, in Sec. 3.2, we provide an overview of the proposed RefSAM model. More details of each component of RefSAM are presented in the remaining sections.
Figure 1: The overall pipeline of RefSAM. It mainly consists of five key components: Visual Encoder of SAM [12], Text Encoder, Cross Modal MLP, Dense Attention, Mask Decoder of SAM. We construct Sparse Embeddings and Dense Embeddings of cross-modal to learn text-visual information and predict masks.
3.1 Preliminaries
Task Definition: Given a video clip $\mathcal{L}$ =$=\{I_{t}\}_{t=1}^{T}$ with $T$ frames and a referring expression $E=\{e_{l}\}_{l=1}^{L}$ with $L$ words, the goal of RVOS is to produce $T$-frame binary segmentation masks $S=\{s_{t}\}_{t=1}^{T}$, $s_{t}\in\mathbb{R}^{H\times W}$ of the referred object.
Architecture of SAM: SAM [12] mainly consists of three components: an image encoder, a prompt encoder, and a mask decoder. The image encoder is a ViT-based [35] backbone for image feature extraction and can be applied prior to prompting the model. The prompt encoder encodes two sets of prompts, i.e., sparse (points and boxes) and dense (masks) prompts, to get interactive positional information and provide them to the mask decoder. The two-layer transformer-based mask decoder leverages the extracted image embedding, prompt embedding, and the learnable output and prompt tokens for final mask prediction. SAM demonstrates strong zero-shot generalization in the segmentation task. However, SAM is not skilled in utilizing text for segmentation and its training is very expensive. For example, training the ViT-H-based SAM model on SA-1B [12] requires 256 GPUs with a large batch size of 256 images.
3.2 Overview of RefSAM’s Architecture
We introduce the RefSAM model to efficiently adapt SAM to the RVOS task and boost the potential segmentation capacity of SAM. As shown in Figure 1, it mainly consists of five key components: Visual Encoder, Text Encoder, Cross-Modal MLP, Dense Attention, and Mask Decoder. Firstly, we use the Visual Encoder of SAM to extract frame features as visual embeddings. Meanwhile, we use the text-to-text model (T5) [53] as the Text Encoder to extract linguistic embeddings. Then, we construct cross-modal Sparse Embeddings and Dense Embeddings to learn text-visual information and predict masks through the Cross-Modal MLP module. Next, RefSAM utilizes the Dense Attetion module to fuse visual embeddings and sparse embeddings in order to obtain dense embeddings. Finally, the Mask Decoder of SAM leverages sparse embeddings, dense embeddings, and visual embeddings for final mask prediction. During training, RefSAM reuses the pre-trained weights of SAM and text encoder T5 while only fine-tuning three modules, i.e., Cross-Modal MLP, Dense Attention, and Mask Decoder, enabling parameter-efficient fine-tuning. During inference, we directly output the mask predictions by selecting the masks with the highest score as the final results.
3.3 Backbone
3.3.1 Visual Encoder
We start by adopting the image encoder of SAM as the visual encoder $\mathcal{E}_{v}$ to extract the visual feature maps for each frame in a video clip, as shown in the light blue part in the top of Figure 1. The image encoder $\mathcal{E}_{v}$ is an MAE [54] pre-trained ViT backbone [35]. Specifically, for each frame $I_{t}$ in the video clip $\mathcal{L}$ =$=\{I_{t}\}_{t=1}^{T}$, the vision encoder $\mathcal{E}_{v}$ is adopted to extract the feature map for this frame, i.e.,
$f_{t}=\mathcal{E}_{v}(I_{t})\in\mathbb{R}^{C_{v}\times H_{0}\times W_{0}},$ (1)
where $f_{t}$ is the corresponding feature map of frame $I_{t}$. By applying the vision encoder $\mathcal{E}_{v}$ on each frame independently, a set of feature maps $\mathcal{F}_{v}=\{f_{t}\}_{t=1}^{T}$ can be obtained for $T$ frames in the video clip. As SAM has strong zero-shot segmentation performance, we freeze the parameters of the image encoder $\mathcal{E}_{v}$ to retain its feature extraction capability in the training process.
3.3.2 Text Encoder
At the same time, given the referring expression with $L$ words, we utilize a large language model text-to-text model (T5) [53] as the text encoder $\mathcal{E}_{t}$ to get the corresponding linguistic embeddings, as shown in the light green part in the bottom of Figure 1. Specifically, given the referring expression $E=\{e_{l}\}_{l=1}^{L}$, these words are firstly tokenized as $T=\{t_{l}\}_{l=1}^{L}$. Then, we put these tokens into the text encoder to obtain the final embeddings. As our text encoder $\mathcal{E}_{t}$ is reused from the T5 model, we extract the feature vector after the last hidden layer of the encoder in the T5 model as the word embedding, i.e.,
$\mathcal{F}_{e}=\mathcal{E}_{t}(E)\in\mathbb{R}^{L\times C_{e}}.$ (2)
Here $\mathcal{F}_{e}$ is a sequence of $C_{e}$-dimensional embeddings of $L$ words, i.e., $\mathcal{F}_{e}=\{f_{i}\}_{i=1}^{L}$, where each word is represented by a $C_{e}$-dimensional embedding. Then the sentence-level embedding $f_{e}^{s}\in\mathbb{R}^{C_{e}}$ can be obtained by simply applying a pooling operation on these word embeddings.
3.4 Cross-modal MLP
Based on the Text Encoder, we get the linguistic embeddings, including word and sentence embeddings. However, there is a significant semantic gap between linguistic embedding space and visual embedding space. Following MiniGPT-4 [55], we design a cross-modal MLP which consists of one hidden layer $L_{s}$ to effectively align the linguistic embedding space and visual embedding space, as shown in the green part in the bottom of Figure 1. Specifically, for each word embedding $f_{i}$ in $\mathcal{F}_{e}$, the sparse embedding can be obtained by adopting the cross-modal MLP $L_{s}$, i.e.,
$f_{i}^{s}=L_{s}(f_{i})\in\mathbb{R}^{C_{v}}.$ (3)
In this way, sparse embeddings for $L$ words and the sentence can be obtained, which can be represented as $\mathcal{F}_{sparse}=\{f_{j}^{s}\}_{j=1}^{L}$ and $f_{e^{\prime}}^{s}\in\mathbb{R}^{C_{v}}$, respectively.
3.5 Dense Attention
To enhance the visual features and provide additional cues for the mask decoder, we introduce the Dense Attention Module to fuse visual and linguistic features, as shown in the orange part in the top middle of Figure 1. Firstly, we concatenate the aligned sentence embeddings $f_{e^{\prime}}^{s}$ and sparse embeddings $\mathcal{F}_{sparse}=\{f_{j}\}_{j=1}^{L}$ together:
$\mathcal{F}_{fix}=(f_{e^{\prime}}^{s}\copyright\mathcal{F}_{sparse})\in\mathbb{R}^{(N+1)\times C_{v}}.$ (4)
In the next step, a similarity calculation between $\mathcal{F}_{fix}$ and visual embeddings $\mathcal{F}_{v}$ is performed to get spatial attention $\mathcal{F}_{sp}$. Then, the dot product is applied between $\mathcal{F}_{sp}$ and $\mathcal{F}_{fix}$ to get spatial language attention, i.e.,
$\mathcal{F}_{sl}=\mathcal{F}_{sp}\cdot\mathcal{F}_{fix}.$ (5)
Then, we concatenate the spatial language attention $\mathcal{F}_{sl}$ and visual embeddings $\mathcal{F}_{v}$ and apply a convolutional layer to reduce the number of channels to match the channel dimension of visual embeddings, i.e.,
$\mathcal{F}_{dense}=Conv(\mathcal{F}_{sl}\copyright\mathcal{F}_{v}).$ (6)
[…]
3.6 Mask Decoder
The mask Decoder of the vanilla SAM leverages sparse embeddings (point and box) from the prompt encoder and dense embeddings (mask) from the SAM predictor to get final predictions. Following this principle, we construct sparse embeddings and dense embeddings with the same shape which encode useful visual and linguistic features from the Cross-Modal MLP and Dense Attention module. Then we put sparse embeddings $\mathcal{F}_{sparse}$ and dense embeddings $\mathcal{F}_{dense}$ together with visual embeddings $\mathcal{F}_{v}$ into the Mask Decoder to get the mask predictions, i.e.,
$M=Decoder(\mathcal{F}_{sparse},\mathcal{F}_{dense},\mathcal{F}_{v}),$ (7)
where $M$ is the output of the Mask Decoder. Finally, the Mask Decoder uses the output $M$ with the highest score as the predicted mask.
4 Experiments
4.1 Experiment Settings
4.1.1 Datasets
We conduct experiments on two challenging referring video object segmentation datasets: Ref-Youtube-VOS [31] and Ref-DAVIS17 [29]. Ref-Youtube-VOS is a large-scale referring video object segmentation dataset, which contains around 15K referring expressions for 3,900+ videos. Each video has pixel-level instance segmentation annotations for every fifth frame. Ref-DAVIS17 is built upon DAVIS17 [30] by providing the language description for each specific object in each video and contains 90 videos. We follow the common practice and use the default split for training and testing.
4.1.2 Evaluation Metrics
The primary evaluation metrics of Ref-Youtube-VOS and Ref-DAVIS17 are the average of the region similarity ($\mathcal{J}$) and the contour accuracy ($\mathcal{F}$), denoted by $\mathcal{J}\&\mathcal{F}$. For Ref-Youtube-VOS, since the annotations of the validation set are not publicly available, we evaluate our method on the official challenge server [^1] https://codalab.lisn.upsaclay.fr/competitions/3282. As for Ref-DAVIS17, it is evaluated using the official evaluation code [^2] https://github.com/davisvideochallenge/davis2017-evaluation.
4.2 Implementation Details
4.2.1 Model Settings
We use the image encoder of SAM, a ViT backbone [35] as our visual encoder. For the text encoder, we use the Hugging Face [56] implementation of T5-3b [53]. The Cross-Modal MLP consists of one hidden layer, which employs the rectified linear unit (ReLU) activation function. The convolution layer we use in Dense Attention applies a set of $512$ $1\times 1$ filters to the input feature maps, resulting in 256 output feature maps. We freeze the parameters of the Visual Encoder and Text Encoder during the entire training stage. We use the last stage features from the Visual Encoder as the input to Dense Attention Module and Mask Decoder. The dimension of the input embeddings (visual embeddings, sparse embeddings, and dense embeddings) to the Mask Decoder is $\mathcal{C}=256$.
4.2.2 Training and Inference Details
Follow [4], our experiments follow the pre-training and fine-tuning process. we perform pre-training on the Ref-COCO dataset [57] for 15 epochs. In practice, we did not reach the optimal number of epochs, but it was sufficient for practical use, demonstrating the efficiency of our method. And then, we fine-tune it on Ref-Youtube-VOS for 4 epochs. The data augmentation includes random horizontal flip, random resize, random crop, and photometric distortion. We employ Dice loss and Focal loss as our primary loss functions. Our model is implemented in PyTorch. Training is conducted on 8 NVIDIA Tesla A100 GPUs.
During inference, the video frames are down-scaled to 360p. We directly output the predicted segmentation masks without using any post-processing tricks.
4.3 Main Results
Table 1: Results on Ref-DAVIS17.
[…]
CMSA [58] 40.2 36.9 43.5
CMSA+RNN [58] 40.2 36.9 43.5
URVOS [31] 51.5 47.3 56.0
ReferFormer [4] 61.1 58.1 64.1
VLT [42] 61.6 58.9 64.3
PerSAM [50]+Grounding Dino [25] 40.3 38.2 42.4
SAM-Track+Grounding Dino [21] 57.1 54.6 59.7
RefSAM 66.1 62.9 69.3
4.3.1 Ref-DAVIS17
[…]
To have a closer look at the effectiveness of our method, we also compare our method with some recent SAM-based methods. Since there are no off-the-shelf SAM-based models for RVOS, existing SAM-based methods, such as SAM-Track [21] and PerSAM [50], can be adapted to RVOS task by providing bounding boxes in the first frame. Specifically, simple SAM-based RVOS baselines can be derived by combining SAM-Track or PerSAM with Grounding Dino [25], which is an object detector that can provide bounding boxes of objects in the first frame with referring expressions. As shown in Table 1, these naive solutions exhibit satisfactory performance, surpassing the CMSA and URVOS methods, yet significantly lagging behind our RefSAM model. Firstly, these methods employ a two-stage pipeline that can lead to a sub-optimal performance in the downstream segmentation task due to inaccurate bounding boxes detected by the object detector in the initial frame. Secondly, the absence of specific fine-tuning on the RVOS dataset poses a challenge, primarily due to the substantial model size and the inherent two-stage design. It is also noteworthy that the involvement of individual models for each sub-task also brings extra difficulties for model deployment. These findings underscore the efficacy of our RefSAM’s end-to-end design and parameter-efficient tuning strategy.
4.3.2 Ref-Youtube-VOS
Table 2: Results on Ref-Youtube-VOS.
[…]
CMSA [58] 34.9 33.3 36.5
CMSA + RNN [58] 36.4 34.8 38.1
URVOS [31] 47.2 45.3 49.2
PMINet [10] 48.2 46.7 49.6
PMINet + CFBI [10] 54.2 53.0 55.5
MTTR [1] 55.3 54.0 56.6
CITD [2] 61.4 60.0 62.7
VLT [42] 63.8 61.9 65.6
ReferFormer [4] 64.9 62.8 67.0
PerSAM [50]+Grounding Dino [25] 30.8 27.1 34.4
SAM-Track+Grounding Dino [21] 52.2 49.9 54.5
RefSAM 55.1 53.9 56.3
[…]
4.4 Ablation Study
We conduct detailed ablation studies on the proposed RefSAM using ViT-B as the backbone, analyzing the impact of the key modules and the influence of hyper-parameters.
4.4.1 Effect of the Key Modules
RefSAM employs a Cross-Modal MLP and Dense Attention for referring video object segmentation. Table 3 summarizes the results of our baseline model when excluding some of the key modules. In each experiment, we first pre-train the models on Ref-COCO for 15 epochs and then fine-tune them on Ref-Youtube-VOS for 4 epochs. We evaluate them on Ref-Davis. Compared to our baseline, without using the Cross-Modal MLP and Dense Attention will lead to a performance drop of 1.56 J&F-Mean and 2.38 J&F-Mean, respectively, validating their importance in aligning and fusing visual and language features.
Table 3: Ablation study of the key components of RefSAM.
[…]
RefSAM 58.5 55.6 61.4
w/o Cross Modal MLP 56.9(-1.6) 55.3 60.6
w/o Dense Attention 56.1(-2.4) 53.4 58.8
4.4.2 Influence of Hyper-parameter Settings
(a) Cross-Modal MLP
(b) Dense Attention
(c) Mask Decoder
Figure 2: The influence of different learning rates for the learnable modules of RefSAM.
(a) Number of hidden layers in Cross-Modal MLP
(b) Linguistic feature in Dense Attention
(c) Model size of Visual Encoder
Figure 3: Ablation study of different module designs.
Figure 2 presents the results of different learning rate settings for the learnable modules of RefSAM, including the Cross-Modal MLP, the Dense Attention, and the Mask Decoder. As can be seen, RefSAM converges faster and better when the learning rate is set to 1e-4 for the Cross-Modal MLP and Dense Attention, and 1e-6 for the Mask Decoder.
Furthermore, we study the influence of the number of hidden layers in the Cross-Model MLP. As shown in Figure 2(a). As can be seen, using a single hidden layer delivers faster and better convergence, which is the default setting. In Figure 2(b), we investigate the choice of linguistic feature in Dense Attention. As can be seen, only using the word feature leads to a slightly better result, which is the default setting.
4.5 Influence of Model Size
We then investigate the influence of model size, specifically the performance of using different sizes of Visual Encoder, including ViT-B, ViT-L, and ViT-H. For each model, we pre-train it on Ref-COCO for 15 epochs, then fine-tune it on Ref-Youtube-VOS for 4 epochs, and evaluate it on Ref-DAVIS17. We do not use the data augmentation described in Sec. 4.2.2. The results are shown in Figure 2(c) and Table 4. The findings indicate that employing a larger backbone for visual feature extraction yields superior performance compared to a smaller one, attributed to the enhanced representation capacity of larger vision transformers. Furthermore, the results exhibit a consistent improvement in performance with increasing model size, highlighting the good scalability of our RefSAM model.
Table 4: Influence of the model size of Visual Encoder in RefSAM.
[…]
ViT-B 58.5 55.6 61.4
ViT-L 62.3 59.0 65.6
ViT-H 63.2 60.0 66.5
(a) A man in a red sweatshirt performing breakdance.
(b) A red bmx bike. A boy wearing a white tshirt.
(c) A horse doing high jumps. A woman riding a horse.
Figure 4: Visualization of RefSAM on Ref-DAVIS17. The text with a specific color corresponds to the object mask of the same color.
(a) A man with a blue belt on the right. A bald man with a black belt in the center.
(b) A black shotting gun. A black man. A rope.
(c) A fat man on the right in a black jacket. A cardboard box held by a man. A man on the left with a beard wearing jeans.
Figure 5: Visualization of different models on Ref-DAVIS17. From left to right: RefSAM, ReferFormer [4], SAM-Track [21] + Ground DINO [25], and PerSAM [50] + Ground DINO. The text with a specific color corresponds to the object mask of the same color.
4.6 Visualization Results
We show the visualization results of our RefSAM model in Figure 4. It can be seen that RefSAM is capable of effectively segmenting and tracking the referred object even in challenging scenarios, such as variations in person poses, and occlusions between instances. Furthermore, we present the results of differnt models in Figure 5. It is clear that our RefSAM demonstrates significantly enhanced cross-modal understanding capability, particularly evident in handling vague language descriptions (Figure 4(a)) and resolving appearance ambiguity among similar objects (Figure 4(c)).
Table 5: Number of learnable parameters of different Models.
Method Visual Encoder #Params (M)
SAM ViT-B 93.7
SAM ViT-L 312.3
SAM ViT-H 641.1
ReferFormer Video-Swin-B 112.9
ReferFormer Swin-L 221.7
RefSAM ViT-B 4.8
RefSAM ViT-L 4.8
RefSAM ViT-H 4.8
Table 6: Inference speed of different models.
Method Backbone FPS
ReferFormer Swin-L 9.1
SAM-Track + Ground DINO ViT-H 1.5
PerSAM + Ground DINO ViT-H 4.0
RefSAM ViT-H 4.6
RefSAM ViT-L 7.0
RefSAM ViT-B 9.2
4.7 Model Complexity Analysis
In this section, we perform an analysis of the complexity of various models. As depicted in Table 5, our model exhibits a notably smaller number of learnable parameters in comparison to all other models, highlighting the inherent nature of parameter-efficient tuning employed during the training of our RefSAM model. Furthermore, Table 6 presents the inference speed of various models. It is apparent that our RefSAM exhibits a marginally slower runtime compared to ReferFormer, yet outperforms SAM-Track and PerSAM in terms of inference speed. When employing a ViT-B as the visual encoder, RefSAM achieves a faster inference speed of 9.2 FPS, outperforming all other methods.
5 Conclusion
This study pioneers the adaptation of the foundational segmentation model, Segment Anything Model, to the referring video object segmentation task. We propose the novel RefSAM model, which incorporates efficient designs to effectively bridge the semantic gap between the visual and language domains, significantly enhancing the cross-modal understanding capability of SAM. By employing a parameter-efficient tuning strategy, RefSAM achieves efficient training through adjustments to a small number of learnable parameters. Extensive experiments on Ref-Youtu-VOS and Ref-DAVIS17 datasets validate the superior effectiveness of our RefSAM model compared to existing methods.
[…]